Write an arithmetic series with seven terms such that the sum of the series is 64
Found 2 solutions by fcabanski, stanbon:
Answer by fcabanski(1391)

The question is mis worded. A sequence is a list of numbers that follows some rule. A series is the sum of a sequence.


An arithmetic sequence has the same number added to each previous number to arrive at the next number. The sum of such a sequence is the sum of the terms.


Pick some number to add to each term. The easiest is 1.


n, n+1, n+2, n+3, n+4, n+5, n+6 is an arithmetic series in which 1 is added to each number to arrive at the next number in the sequence.


The sum is n + n+1 + n+2 + n+3 + n+4 + n+5 + n+6 and it must = 64.


n + n+1 + n+2 + n+3 + n+4 + n+5 + n+6 = 64


7n + 21=64


7n = 43
n = 43/7


The sequence is 43/7, 50/7, 57/7, 64/7, 71/7, 78/7, 85/7


Answer by stanbon(75887)

Write an arithmetic series with seven terms such that
the sum of the series is 64.
Let the 1st term be "a";
Then the 7th term is a + 6d
-----------------------------------
S(7) = (7/2)(a+(a+6d)) = 64
----
7(2a+6d) = 128
7(a+3d) = 64
----
7a+21d = 64
a + 3d = 9.1429
----
Let a = 0.1426
Then d = 3
----
0.1429,3.1429,6.1429,9.1429,12.1429,15.1429,18.1429
==================
Cheers,
Stan H.
====================